OpenStudy (anonymous):

A truck weighs 4500 kg. The truck is packed with bales of wool, each weighing 150 kg. Part A: Write a function to show the total weight of the truck f(w) when it is loaded with n bales of wool. (4 points) Part B: What is the total weight of the truck when it is packed with 15 bales of wool? (4 points) Part C: If a carton weighing 20 kg is also loaded on the same truck, what would be the weight of the truck when it is loaded with 10 bales of wool? (2 points) I don't know how to answer this can anyone help me with the steps?

OpenStudy (triciaal):

total weight is the weight of the truck plus the weight of all the wool. the weight of all the wool is the number of bales times the weight of each bale.

OpenStudy (anonymous):

but im not sure how to write that as a function is it 4500kg=150kg(x)

OpenStudy (triciaal):

now can you write the function if you let the number of bales = x and the total weight = y

OpenStudy (triciaal):

remember y = mx + c c would be the weight of the truck by itself

OpenStudy (anonymous):

so my function is wrong?

OpenStudy (triciaal):

yes

OpenStudy (triciaal):

read what i wrote to start

OpenStudy (triciaal):

total weight is the weight of the truck plus the weight of all the wool.

OpenStudy (anonymous):

should it be 150kg(x)=4500kg?

OpenStudy (anonymous):

4500+150kg=x?

OpenStudy (triciaal):

no the weight varies with the number of bales of wool the weight of the truck is the toal weight when there is no wool

OpenStudy (triciaal):

y = mx + c total weight = weight of one bale times the number of bales added to the weight of the truck

OpenStudy (triciaal):

f(x) = 150x + 4500

OpenStudy (anonymous):

oh ok

OpenStudy (triciaal):

do you see that now?

OpenStudy (anonymous):

yes thank you so for part b you would write f(x)=150(15)+4500 right

OpenStudy (triciaal):

total weight of the truck f(w)

OpenStudy (triciaal):

part b yes

OpenStudy (anonymous):

thanks for part c I don't know what to do

OpenStudy (triciaal):

yes you do! you have a truck with 10 bales of wool and then I gave you your lunch bag?

OpenStudy (anonymous):

so u add the 10 and 20 and get 30 then you say f(x)= 150(30)+4500

OpenStudy (triciaal):

150 is the weight of 1 bale of wool and you have 10 plus the weight of the truck plus the carton sorry I was hoping you would know to add your lunch bag the carton so add 20

OpenStudy (triciaal):

total weight = truck + wool + carton = + 150(10) + 4500 + 20

OpenStudy (anonymous):

oh so its 1500+4520=6020

OpenStudy (triciaal):

got your 2 points!

OpenStudy (anonymous):

thanks!

OpenStudy (triciaal):

no problem, you are welcome